A cyclist traveled at a rate of 32 mph to vist a nearby town. The cyclist averaged 10 mph on the return trip. If the round trip took 6.3 hours, find the distance to the town.


Answer by nerdybill(7384)

A cyclist traveled at a rate of 32 mph to vist a nearby town. The cyclist averaged 10 mph on the return trip. If the round trip took 6.3 hours, find the distance to the town.
.
Let x = time it took to get to the town
then
6.3-x = time it took to return
.
Applying the distance formula: d = rt
.
32x = 10(6.3-x)
32x = 63-10x
42x = 63
x = 63/42
x = 1.5 hours
.
Distance to town:
1.5(32) = 48 miles